OpenStudy (anonymous):

What is the product in simplest form,state any restrictions. x^2+5x+6/x+4 * x^2+x-12/x^2+x-2 Help please! I'm drowning in a sea of confusion

OpenStudy (luigi0210):

This? \(\Large \frac{x^2+5x+6}{x+4} ~*~ \frac{x^2+x-12}{x^2+x-2} \)

OpenStudy (anonymous):

Yes ! exactly that

OpenStudy (luigi0210):

Alright, first thing you'll need to do is factor, can you do that?

OpenStudy (anonymous):

Which part do i factor first?

OpenStudy (luigi0210):

factor \(x^2+5x+6 \), what do you get?

OpenStudy (anonymous):

(x+2) (x+4) I'm not sure how to factor the 5x. I know you need to add to get both 5x and 6

OpenStudy (anonymous):

(x+2) (x+3)

OpenStudy (luigi0210):

There ya go! :) Now, factor \(x^2+x-12 \)

OpenStudy (anonymous):

(x+6) (x-2) that gives the -12 but not the x in the middle, i just need another moment

OpenStudy (luigi0210):

Take your time (:

OpenStudy (anonymous):

(x-4) (x+3)

OpenStudy (anonymous):

switch the positive and negative sign though so (x+4) (x-3)

OpenStudy (luigi0210):

Close enough :P That gives us \(x^2-x-12 \) so just change the signs and you're good I'll do the last one for ya if you don't mind :P After factoring it all, we get this: \(\Large \frac{(x+2)(x+3)}{(x+4)}*\frac{(x+4)(x-3)}{(x+2)(x-1) } \) Now we need to cancel things; notice anything that simplies?

OpenStudy (anonymous):

Could you cross out the (x+2) ?

OpenStudy (luigi0210):

Yes we could! \(\Large \frac{\cancel{{(x+2)}}(x+3)}{(x+4)}*\frac{(x+4)(x-3)}{\cancel{(x+2)}(x-1) } \) Anything else?

OpenStudy (anonymous):

those damn (x+4)'s xD

OpenStudy (luigi0210):

Yup \(\Large \frac{\cancel{{(x+2)}}(x+3)}{\cancel{(x+4)} }*\frac{ \cancel{(x+4)} (x-3)}{\cancel{(x+2)}(x-1) } \) Which simplies to just: \(\Large \frac{(x+3)(x-3)}{x-1} \) Now just gotta find any restrictions (:

OpenStudy (anonymous):

3, -3 and 1? or do you not include the factored expression on top

OpenStudy (luigi0210):

Restrictions are usually just in the denominator, so \(\ x \neq 1 \)

OpenStudy (anonymous):

Alright, thank for taking so long to explain this too me CX I dig your hair too! Enjoy the rest of the day

OpenStudy (luigi0210):

Haha, no problem, and thanks, you too :D